Price of trout (X) per pound is $2. Price of salmon (Y) per pound is $10. What is oC of buying a pound of salmon i.e. how much trout do you give up if you buy a pound of salmon? What will be the relative price of buying a pound of salmon?
Answer #1

Answer :

Given price of trout =$2

Price of salmon=$10

OC = price of salmon/price of trout = 10/2=5 pound of Trout

Relative price of salmon=10/2=$5